# Decomposing Representation Space into Interpretable Subspaces with Unsupervised Learning

### v1

@@ F.1 Why is InversionView More Suitable in Subspaces
Compared to whole representation space, we argue that InversionView [Huang et al., 2024b] is more suitable to be applied to subspaces. For example, subspace 2 and subspace 7 in Fig. 2 encode current token and current position respectively. If we look for the preimage in the joint space of subspace 2 and 7, the inputs will need to have the same token at the same position to be inside the preimage. Thus when the query activation contains multiple pieces of information, or encode multiple aspects of the input, the samples in preimage should be the same in all these aspects, which makes it hard to find such samples. Even if we can find enough, they might simply have the same recent tokens, which does not give much insights.
@@
F.2 On the Reliability of Subspace Activation Patching
@@
Figure 31: An example illustrating subspace activation patching illusion. Fig. 1 from [Makelov et al., 2023]. The 2D space is partitioned into two 1D space (red arrrows). Green arrows represent activations in the whole space. Subspace patching modify their projection along the red $\boldsymbol{v}$ direction, resulting a new activation (blue arrow). In this setting, suppose x-axis has variance but no downstream causal effect, while y-axis has no variance when running model on normal data but has causal effect if there is any nonzero value. In this setting, resulting new activation (blue arrow) is out-of-distribution and will have causal effect, suggesting $\boldsymbol{v}$ is a meaningful direction. However, this 2D space does not play a meaningful role in model’s computation (suppose this 2D space is a subspace of a bigger space). In our case this does not raise a severe concern. When applying NDM in this 2D space, partitioning like red arrows in this figure is not likely to happen, because the resulting two 1D subspaces have very high MI. Partitioning like the x and y-axis is the optimal choice, thus resulting two unimportant 1D space.
@@
Our quantitative experiments on language models relies on subspace activation patching. Makelov et al. [2023] found cases where subspace activation patching produces illusory conclusion. Subspace intervention might change the model’s output by activating a dormant pathway which does not have effect in normal computation, as illustrated in Fig. 31. Here we argue that our experiments are still reliable because of following reasons: (1) Our method finds subspace partition that reduces MI or dependency between subspaces, instead of partition that maximize patching effect. The partition shown in Fig. 31 that causes illusion has very high MI between subspaces, thus our method would not give such partition. In other words, we apply subspace patching on subspaces found by NDM, because of relatively good independency, so activations from different subspaces can be composed more or less freely, without causing out-of-distribution activations. (2) Makelov et al. [2023] recommend doing subspace patching in activation bottlenecks, especially residual stream, to avoid such illusion. This is exactly what we did in experiments (3) We also provide validation by examining the subspaces qualitatively.
F.3 Possible Alternative Approaches
@@
    partition the space into two subspaces. We search over n possible configurations (for example, when d=100, we can have [$d_{1}=25,d_{2}=75$], [$d_{1}=50,d_{2}=50$], [$d_{1}=75,d_{2}=25$]), each configuration has a dedicated orthogonal matrix. It turns out that initialization matters a lot and standard basis is kind of special (Fig. 14), though they are symmetric, empirically we found [$d_{1}=25,d_{2}=75$] and [$d_{1}=75,d_{2}=25$] produces quite different results. After certain amount of steps, measure (normalized) MI for each configuration, select the one with the lowest MI.
@@
    Repeat step 1 for each subspaces already obtained, while restricting the rotation inside the already obtained subspace. For example, assume we have learned a matrix $\mathbf{R}^{\text{prev}}$ and configuration [$d_{1}^{\text{prev}}=75,d_{2}^{\text{prev}}=25$], we keep (previous) subspace 1 and 2 fixed. For subspace 1, we search over configurations [$d_{1}=18,d_{2}=57$], [$d_{1}=36,d_{2}=39$], [$d_{1}=54,d_{2}=21$], each associated with a separate $\mathbf{R}_{1},\mathbf{R}_{2},\mathbf{R}_{3}\in\mathbb{R}^{75\times 75}$. Similarly do this for subspace 2. Assume in this iteration, after training [$d_{1}=36,d_{2}=39$] has lowest MI for subspace 1, we then merge its orthogonal matrix $\mathbf{R}_{2}$ into $\mathbf{R}^{\text{prev}}$ by multiplying it with corresponding rows in $\mathbf{R}^{\text{prev}}$: $\mathbf{R}^{\text{prev}}_{[0:75]}\leftarrow\mathbf{R}_{2}\mathbf{R}^{\text{prev}}_{[0:75]}$. On the other hand, suppose none of the configuration for subspace 2 is lower than threshold, then we do not update that subspace. After this iteration, we have a new $\mathbf{R}^{\text{prev}}$, and new configuration [$d_{1}^{\text{prev}}=36,d_{2}^{\text{prev}}=39,d_{3}^{\text{prev}}=25$]. Then keep repeating this loop until no subspace is separable (all configurations of all subspaces have MI greater than the threshold), or a pre-defined maximum iteration is reached.
@@
$\text{Var}_{s}$ (×$\times 10^{3}$) 0.4 1.0 0.5 0.4 0.8 2.7 8.3 0.6 0.4 0.6 0.6 5.8
$\Delta P_{s}$ (×$\times 10^{-2}$) 0.3 1.3 1.1 0.3 0.3 2.8 2.3 0.6 0.4 1.9 0.6 40.7
$\frac{\Delta P_{s}}{d_{s}}$ (×$\times 10^{-4}$) 0.9 2.5 2.0 0.8 2.1 1.9 2.2 1.6 2.3 3.5 1.1 25.1
$\frac{\Delta P_{s}}{\text{Var}_{s}}$ (×$\times 10^{-5}$) 0.9 1.3 2.2 0.7 0.3 1.1 0.3 0.9 1.1 3.2 1.0 7.0
@@
Mutual information can also be estimated by neural networks [Belghazi et al., 2018]. The Mutual Information Neural Estimator (MINE) is as follows:
Let =$\mathcal{F}=\{T_{\theta}\}_{\theta\in\Theta}$ be the set of functions parametrized by a neural network. MINE is defined as (we simplify the notation to avoid some details, please refer to [Belghazi et al., 2018] for the true definition):
@@
The intuition is that, if X and Z have high mutual information, knowing one can predict the other, the neural network can easily discriminate whether $\boldsymbol{x}$ and $\boldsymbol{z}$ are projected from the same activation, thus output high value when they “match” and low value when they do not, resulting high estimate of $I(X;Z)$. $\mathbf{R}$ rotates and reflects the space such that two subspace are as independent as possible, such that the neural network cannot easily discriminate, resulting low MI estimate. This is similar to Generative Adversarial Networks [Goodfellow et al., 2020].
@@
Appendix G Compute Usage

### v2

@@ G.1 Why is InversionView More Suitable in Subspaces
Compared to whole representation space, we argue that InversionView (Huang et al., 2024b) is more suitable to be applied to subspaces. For example, subspace 2 and subspace 7 in Fig. 3 encode current token and current position respectively. If we look for the preimage in the joint space of subspace 2 and 7, the inputs will need to have the same token at the same position to be inside the preimage. Thus when the query activation contains multiple pieces of information, or encode multiple aspects of the input, the samples in preimage should be the same in all these aspects, which makes it hard to find such samples. Even if we can find enough, they might simply have the same recent tokens, which does not give much insights.
@@
G.2 On the Reliability of Subspace Activation Patching
@@
Figure 33: An example illustrating subspace activation patching illusion. Fig. 1 from (Makelov et al., 2023). The 2D space is partitioned into two 1D space (red arrrows). Green arrows represent activations in the whole space. Subspace patching modify their projection along the red $\boldsymbol{v}$ direction, resulting a new activation (blue arrow). In this setting, suppose x-axis has variance but no downstream causal effect, while y-axis has no variance when running model on normal data but has causal effect if there is any nonzero value. In this setting, resulting new activation (blue arrow) is out-of-distribution and will have causal effect, suggesting $\boldsymbol{v}$ is a meaningful direction. However, this 2D space does not play a meaningful role in model’s computation (suppose this 2D space is a subspace of a bigger space). In our case this does not raise a severe concern. When applying NDM in this 2D space, partitioning like red arrows in this figure is not likely to happen, because the resulting two 1D subspaces have very high MI. Partitioning like the x and y-axis is the optimal choice, thus resulting two unimportant 1D space.
@@
Our quantitative experiments on language models relies on subspace activation patching. (Makelov et al., 2023) found cases where subspace activation patching produces illusory conclusion. Subspace intervention might change the model’s output by activating a dormant pathway which does not have effect in normal computation, as illustrated in Fig. 33. Here we argue that our experiments are still reliable because of following reasons: (1) Our method finds subspace partition that reduces MI or dependency between subspaces, instead of partition that maximize patching effect. The partition shown in Fig. 33 that causes illusion has very high MI between subspaces, thus our method would not give such partition. In other words, we apply subspace patching on subspaces found by NDM, because of relatively good independency, so activations from different subspaces can be composed more or less freely, without causing out-of-distribution activations. (2) (Makelov et al., 2023) recommend doing subspace patching in activation bottlenecks, especially residual stream, to avoid such illusion. This is exactly what we did in experiments (3) We also provide validation by examining the subspaces qualitatively.
G.3 About Future Directions
So far we have provided various evidence supporting our claim that we can decompose representation space into smaller and meaningful subspaces, without any human supervision. These subspaces, being orthogonal to each other, can serve as new basic units or mediators for mechanistic interpretability. Compared to neurons, they capture the distributed nature of neural representations. Compared to sparse features, they provide more structure. Each subspace can be thought of as a group of infinitely many features sharing a certain higher-level concept, by interpreting some of them we can understand the whole group. Moreover, it enables us to understand the model by directly analyzing the model’s weight matrices. For example, we can understand an attention head by checking which subspace its query, key, and value matrices read from and which subspace it writes to (Elhage et al., 2021). By doing so, we build connections between subspaces across layers, resulting in subspace circuits. We believe subspaces could be more “fundamental” than model components as they are designed to be as independent as possible. Moreover, subspace circuits would, like real transformers, operate in the same manner over the whole input sequence (uniform across token positions) as well as over different input sequences. Such circuits could be predecessors of source-code-like explanation of the model, since variables in RASP language (Weiss et al., 2021) can fundamentally be treated as subspaces of residual stream. To support this ambitious plan, the current version of subspace partition method might not be good enough. We believe there is plenty of room for further improvement. For example, allowing more flexible searching over dimension configurations (rather than having all dimensions as multiples of a fixed number). We also think a hierarchical structure of subspaces could also be helpful. For example, subspace A and B are very independent (e.g., current token and position), and A can be further partitioned into A1 and A2, which are less independent (e.g., the plain token and retrieved knowledge about it). More importantly, there could be potentially different ways achieving the same goal (e.g., by minimax approaches, see App. G.4). We also call for efforts from industry as this unsupervised method should benefit from more data and compute.
G.4 Possible Alternative Approaches
@@
    partition the space into two subspaces. We search over n possible configurations (for example, when d=100, we can have $d_{1}=25,d_{2}=75$], $d_{1}=50,d_{2}=50$], $d_{1}=75,d_{2}=25$]), each configuration has a dedicated orthogonal matrix. It turns out that initialization matters a lot and standard basis is kind of special (Fig. 15), though they are symmetric, empirically we found $d_{1}=25,d_{2}=75$] and $d_{1}=75,d_{2}=25$] produces quite different results. After certain amount of steps, measure (normalized) MI for each configuration, select the one with the lowest MI.
@@
    Repeat step 1 for each subspaces already obtained, while restricting the rotation inside the already obtained subspace. For example, assume we have learned a matrix $\mathbf{R}^{\text{prev}}$ and configuration $d_{1}^{\text{prev}}=75,d_{2}^{\text{prev}}=25$], we keep (previous) subspace 1 and 2 fixed. For subspace 1, we search over configurations $d_{1}=18,d_{2}=57$], $d_{1}=36,d_{2}=39$], $d_{1}=54,d_{2}=21$], each associated with a separate $\mathbf{R}_{1},\mathbf{R}_{2},\mathbf{R}_{3}\in\mathbb{R}^{75\times 75}$. Similarly do this for subspace 2. Assume in this iteration, after training $d_{1}=36,d_{2}=39$] has lowest MI for subspace 1, we then merge its orthogonal matrix $\mathbf{R}_{2}$ into $\mathbf{R}^{\text{prev}}$ by multiplying it with corresponding rows in $\mathbf{R}^{\text{prev}}$: $\mathbf{R}^{\text{prev}}_{[0:75]}\leftarrow\mathbf{R}_{2}\mathbf{R}^{\text{prev}}_{[0:75]}$. On the other hand, suppose none of the configuration for subspace 2 is lower than threshold, then we do not update that subspace. After this iteration, we have a new $\mathbf{R}^{\text{prev}}$, and new configuration $d_{1}^{\text{prev}}=36,d_{2}^{\text{prev}}=39,d_{3}^{\text{prev}}=25$]. Then keep repeating this loop until no subspace is separable (all configurations of all subspaces have MI greater than the threshold), or a pre-defined maximum iteration is reached.
@@
$\text{Var}_{s}$ ($\times 10^{3}$) 0.4 1.0 0.5 0.4 0.8 2.7 8.3 0.6 0.4 0.6 0.6 5.8
$\Delta P_{s}$ ($\times 10^{-2}$) 0.3 1.3 1.1 0.3 0.3 2.8 2.3 0.6 0.4 1.9 0.6 40.7
$\frac{\Delta P_{s}}{d_{s}}$ ($\times 10^{-4}$) 0.9 2.5 2.0 0.8 2.1 1.9 2.2 1.6 2.3 3.5 1.1 25.1
$\frac{\Delta P_{s}}{\text{Var}_{s}}$ ($\times 10^{-5}$) 0.9 1.3 2.2 0.7 0.3 1.1 0.3 0.9 1.1 3.2 1.0 7.0
@@
Mutual information can also be estimated by neural networks (Belghazi et al., 2018). The Mutual Information Neural Estimator (MINE) is as follows:
Let $\mathcal{F}=\{T_{\theta}\}_{\theta\in\Theta}$ be the set of functions parametrized by a neural network. MINE is defined as (we simplify the notation to avoid some details, please refer to (Belghazi et al., 2018) for the true definition):
@@
The intuition is that, if X and Z have high mutual information, knowing one can predict the other, the neural network can easily discriminate whether $\boldsymbol{x}$ and $\boldsymbol{z}$ are projected from the same activation, thus output high value when they “match” and low value when they do not, resulting high estimate of $I(X;Z)$. $\mathbf{R}$ rotates and reflects the space such that two subspace are as independent as possible, such that the neural network cannot easily discriminate, resulting low MI estimate. This is similar to Generative Adversarial Networks (Goodfellow et al., 2020).
@@
Appendix H Compute Usage
@@
Appendix I LLMs Usage
We used LLMs to polish the writing and rephrase some statements to improve the readability, while ensuring that the content continues to convey our intended meaning accurately. We also used them to identify related work and to generate code for producing figures.